OpenStudy (anonymous):

If the graph of the function defined by f(x)=-4x^2-8 is translated vertically upward by 5 units, it becomes the graph of a function g . Find the expression for . g(x)

OpenStudy (anonymous):

All you have to do is make every value of f(x) 5 more. So add 5 to f(x) g(x) = f(x) + 5 = (-4x^2 - 8) + 5 g(x) = -4x^2 -3 Hope that helps!

OpenStudy (anonymous):

thank you! was able to get the whole section on those done with that advice ty!